Could someone please show me how to solve by substitution method?
8x+8y=-8
-2x+y=14

Answer by rfer(16322)

-2x+y=14 y=2x+14
8x+8y=-8
8x+8(2x+14)=-8
8x+16x+112=-8
24x=-120
x=-5
----------------
8*-5+8y=-8
-40+8y=-8
8y=32
y=4
--------------
proof
-2*-5+4=14